OpenStudy (anonymous):

What is the value of x in the equation 1/5x -2/3 x – y = 30, when y = 15?

OpenStudy (anonymous):

Is this it? \(\dfrac{1}{5}x - \dfrac{2}{3}x - y = 30\)